# SemViQA: A Semantic Question Answering System for Vietnamese Information Fact-Checking

### v2

@@ 4.2 Experimental Setup
We optimized our model through extensive experiments on an NVIDIA A100 GPU, fine-tuning key hyperparameters while maintaining consistent settings across all runs. The final hyperparameters, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC. Details are provided in Appendix B. For evaluation, all methods were tested on an NVIDIA T4 instance in Kaggle to ensure fairness.
Additionally, the large language model was fine-tuned on a distributed setup with A100 GPUs, utilizing a structured prompt reformulation process. The dataset, originally in raw text and labels, was transformed into a prompt-based format optimized for LLM fine-tuning. This ensured effective task-specific learning while aligning with pre-trained knowledge. Further details on training setup, prompt design, and preprocessing are in Appendix B.
@@ 4.3.1 Performance Comparison
One major limitation of conventional Question Answering (QA) models in fact verification tasks is their inability to process long-context claims due to token length constraints. Transformer-based models such as ViMRClarge, InfoXLMlarge Chi et al. (2021), and XLM-Rlarge Conneau et al. (2020), Ernie-Mlarge Ouyang et al. (2021) are restricted to a maximum input length of 512 tokens. However, real-world fact-checking datasets, such as ISE-DSC01, often contain contexts exceeding 4800 tokens, causing a significant drop in performance for QA-based approaches due to their inability to handle long-token sequences. This limitation prevents these models from effectively leveraging the full evidence available, leading to incomplete or inaccurate fact verification.
To address this issue, our proposed SemViQA model incorporates a retrieval-based approach capable of handling long-token sequences efficiently, as illustrated in Figure 3. On ISE-DSC01, where long contexts are essential for verification, SemViQA significantly outperforms traditional QA models by effectively capturing and utilizing extended contextual information. This demonstrates that the long-token issue is a critical bottleneck in fact verification tasks.
In contrast, on the ViWikiFC dataset, where the token length is closer to 512 tokens, QA-based models achieve competitive performance, further confirming that the long-token issue is a major factor in performance degradation. However, when the long-token limitation does not exist on the ViWikiFC dataset, integrating our SER approach into a QA-based model proves to be an effective solution. By leveraging SER, we achieve approximately 1.86% higher ER Accuracy on ViWikiFC, demonstrating its efficiency in optimizing fact verification when token constraints are minimized.
These results highlight that long-token constraints are a fundamental challenge in fact verification. Our approach not only mitigates this issue through retrieval-based processing in SemViQA but also enhances QA models with QACT when long-token constraints are less severe, ensuring improved performance across diverse datasets.
@@
One of the key advancements of SemViQA over other methods is its significant reduction in inference time without compromising accuracy. Key observations include:
The average inference time of SemViQA on ISE-DSC01 is 5200s, whereas large LLM-based models such as Qwen2.5-3B-Instruct Qwen et al. (2024) require over 31,000s, making SemViQA at least 6 times faster than these large language model approaches.
Compared to ViMRClarge333https://huggingface.co/nguyenvulebinh/vi-mrc-large (9800s on ISE-DSC01), SemViQA reduces inference time by nearly 50% while maintaining higher performance in both Strict Accuracy and Veracity Classification Accuracy.
Methods based on SBERT Reimers and Gurevych (2019b) or BM25 have lower inference times but fail to maintain high accuracy, particularly when dealing with multi-step reasoning tasks. SemViQA achieves a better balance between classification performance and processing speed, ensuring practical deployment in real-world environments.
SemViQA Faster: We introduce an optimized version of SemViQA that achieves up to 7× faster inference while maintaining high accuracy. As shown in Figure 4, the inference speed of SER-Faster closely matches that of traditional methods while still outperforming competing solutions in both efficiency and accuracy. To achieve this speed improvement, we process all subcontexts within a sample as a batch, enabling batch inference (see Section 3.1 for details on subcontexts). While this approach results in a slight performance drop compared to the standard SemViQA model, the trade-off is minimal considering the significant reduction in processing time. This efficiency enhancement makes SemViQA Faster highly suitable for real-world deployment, allowing seamless integration into large-scale fact-checking applications.
4.3.2 Analysis of Confidence Threshold in SemViQA
The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process. Analysis from Figure 5 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%). However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively. Setting an optimal threshold in the range of 0.4 – 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
@@ 4.3.3 Comparison with other results in the competition
| DS@UIT Dynasty | 78.05 | 84.76 | 80.13 |
| URA_FNU | 77.87 | 83.71 | 79.96 |
| Plain Sailing | 77.09 | 82.33 | 78.31 |
| ViNSV | 76.33 | 81.67 | 78.11 |
@@ 5 Conclusion and Future Works
Appendix A Strict Accuracy in Fact-Checking
@@
Appendix B Hyperparameter and LLM Training Configuration
@@
In this section, we present the detailed hyperparameter settings and training configurations for both our SemViQA models and the Large Language Model (LLM) fine-tuning process. Table 4 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering Token Classifier (QATC), and LLM fine-tuning.
@@
We fine-tune a Large Language Model (LLM) using a restructured version of the original datasets, ViWikiFC and ISE-DSC01, as detailed in Figure 5. These datasets have been carefully adapted for training to improve performance and ensure compatibility with our model. For training, we utilize the official Qwen LLM implementation from the QwenLM repository444https://github.com/QwenLM/Qwen. Our training setup follows the full configuration outlined in Table 4, ensuring optimal efficiency and alignment with best practices.
[backgroundcolor=gray!5, linewidth=0.75pt, roundcorner=5pt]
Question: You are tasked with verifying the correctness of the following statement.
@@
- Remember, the evidence must be a full sentence, not part of a sentence or less than one sentence.
Given a claim and context as follows:
Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day. For dinner, Ivy Chen usually eats early to ensure her body has enough time to digest food, metabolize energy, and avoid putting pressure on the stomach and other organs. A recent study published in *Frontiers in Nutrition* suggests that eating dinner earlier can lead to a longer lifespan, with the ideal time being 7 PM. If this is not possible, experts recommend having the last meal of the day 2-3 hours before bedtime. Drinking ginger tea: To keep her body warm, promote blood circulation, and enhance circulation, Ivy Chen drinks ginger tea daily. Her ginger tea is typically made with ground ginger, black tea, turmeric powder, and brown sugar. This drink is a natural remedy that not only boosts the immune system and reduces inflammation but also fights oxidation, supports weight loss, improves skin health, and helps maintain a youthful look. Regular exercise: Ivy Chen is a fitness enthusiast who loves physical activities and exercises daily, even during pregnancy. The Taiwanese actress shared that if she is not busy with work, she runs for at least 30 minutes every day. Even when traveling abroad, she maintains her running habit. A recent study published in *Progress in Cardiovascular Disease* found that regular runners live three years longer than non-runners. Running significantly helps with weight loss, maintaining a balanced physique, toning muscles, relaxing the mind, and benefiting heart health. Besides running, Ivy Chen also swims, practices yoga, and hikes to maintain physical fitness and endurance. Skincare: Regarding her skincare routine, the actress emphasized the importance of hydration. The Taiwanese beauty revealed that she always carries a facial mist to ensure her skin stays hydrated while outdoors.
Claim: Even when traveling abroad, Ivy Chen maintains her running habit.
Answer: This claim is classified as SUPPORTED. The evidence is: Even when traveling abroad, she maintains her running habit.
@@
Appendix C Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection

### v3

@@ 4.2 Experimental Setup
We conducted extensive experiments on NVIDIA A100 GPUs, fine-tuning key hyperparameters while keeping consistent settings across runs. The final configuration, selected via rigorous validation, improved both accuracy and strict accuracy on ISE-DSC01 and ViWikiFC. Full details are provided in Appendix D. For fair evaluation, all methods were tested on a Kaggle instance with an NVIDIA T4 GPU.
The large language model was fine-tuned in a distributed A100 setup using a structured prompt-based reformulation. Raw data were converted into prompt format to align with LLM training objectives and maximize task-specific performance. Training setup, prompt design, and preprocessing are also detailed in Appendix D.
@@ 4.3.1 Performance Comparison
A major limitation of conventional Question Answering (QA) models in fact verification is their inability to process long-context claims due to the 512-token input limit of transformer-based models such as ViMRClarge333https://huggingface.co/nguyenvulebinh/vi-mrc-large, InfoXLMlarge Chi et al. (2021), XLM-Rlarge Conneau et al. (2020), and Ernie-Mlarge Ouyang et al. (2021). Real-world datasets like ISE-DSC01 often contain contexts exceeding 4800 tokens, severely degrading QA-based performance by limiting access to full evidence. To overcome this, SemViQA employs an efficient retrieval-based strategy (see Section 3.1.1) that handles long-token sequences effectively. On ISE-DSC01, SemViQA outperforms traditional QA models by fully leveraging extended contexts, confirming that the long-token constraint is a critical bottleneck. Conversely, on ViWikiFC, where contexts average around 512 tokens, QA models perform competitively. Yet, even in this setting, integrating our Semantic-based Evidence Retrieval (SER) yields a 1.86% improvement in evidence retrieval accuracy, demonstrating the versatility and efficiency of our approach. These findings emphasize that long-token limitations significantly hinder fact verification, and SemViQA successfully mitigates this issue while enhancing QA models across varying dataset conditions.
@@
SemViQA significantly reduces inference time while maintaining high accuracy, making it highly practical for real-world applications. Key highlights include:
On ISE-DSC01, SemViQA averages 5200s per run, over 6 times faster than large LLM-based models like Qwen2.5-3B-Instruct Qwen et al. (2024), which require over 31,000s.
Compared to ViMRClarge (9800s), SemViQA halves inference time while achieving superior Strict Accuracy and Veracity Classification Accuracy.
Although BM25 and SBERT Reimers and Gurevych (2019b) are faster, they struggle with complex, multi-step reasoning where SemViQA maintains a strong balance between speed and accuracy.
SemViQA Faster: We further introduce SemViQA Faster, which accelerates inference by up to 7 through batch processing of subcontexts (see Section 3.1.1). As shown in Figure 4, this variant achieves inference speeds comparable to traditional methods while retaining competitive accuracy. The minor performance trade-off is acceptable given the substantial time savings, making SemViQA Faster ideal for scalable, real-world fact-checking systems.
@@ 4.3.2 Comparison with other results in the competition
| DS@UIT Dynasty444https://github.com/minhquan6203/Verdict-Classification-for-Fact-Checking-at-DSC-2023 | 78.05 | 84.76 | 80.13 |
| URA_FNU555https://github.com/virrosluo/URA_UIT_Data_Science_Challenge | 77.87 | 83.71 | 79.96 |
| ViNSV Tran et al. (2024b) | 76.33 | 81.67 | 78.11 |
| Tran et al. (2024a) | 75.11 | 82.30 | 76.82 |
@@ 5 Conclusion and Future Works
Appendix A Comprehensive Data Processing
A.1 Dataset Statistics
This section presents the detailed statistics of the two datasets used in our experiments: ISE-DSC01 and ViWikiFC. The table below summarizes the number of samples in the training, development, and test sets for both corpora. These statistics help contextualize the scale and evaluation coverage of our methods.
| ISE-DSC01 | ViWikiFC | |
| Train | 37,967 | 16,738 |
| Dev | 4,794 | 2,090 |
| Test | 5,396 | 2,091 |
A.2 Data Analysis on Context Lengths
Figure 5 illustrates the distribution of token lengths across input contexts in the ViWikiFC and ISE-DSC01 datasets. As shown, many samples significantly exceed the 512-token input limitation of standard Transformer models. The ISE-DSC01 dataset, in particular, contains several contexts with over 4,000 tokens. This analysis highlights the necessity for effective context segmentation strategies to ensure full coverage of relevant evidence while maintaining compatibility with model constraints.
Figure 3 illustrates our strategy for handling long input contexts. First, the context is segmented into individual sentences. Next, sentences are sequentially aggregated into subcontexts until reaching approximately 400 tokens. Each completed subcontext is then passed to the QATC model to identify potential evidence. The next subcontext begins from the subsequent sentence, continuing until all sentences have been processed. However, processing subcontexts sequentially can be time-consuming. Therefore, we developed SER Faster, which batches subcontexts and processes them in parallel, significantly accelerating the retrieval process.
Appendix B Strict Accuracy in Fact-Checking
@@
Appendix C Ablation Study
C.1 Ablation Experiments
| Configuration | ViWikiFC | ISE-DSC01 | |||||
| ER Model | VC Model | Strict Acc | VC Acc | ER Acc | Strict Acc | VC Acc | ER Acc |
| SemViQA (Full) | |||||||
| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 75.13 | 79.60 | 76.87 |
| XLM-R | 80.82 | 83.88 | 95.31 | 76.74 | 81.71 | 78.95 | |
| Ernie-M | 80.06 | 83.17 | 95.31 | 78.97 | 82.49 | 80.91 | |
| w/o Binary Classification (One-step VC) | |||||||
| InfoXLM | InfoXLM | 79.63 | 82.88 | 95.31 | 73.87 | 78.35 | 76.89 |
| XLM-R | 80.73 | 83.69 | 95.31 | 75.96 | 80.80 | 78.97 | |
| Ernie-M | 79.91 | 83.07 | 95.31 | 78.47 | 81.89 | 80.93 | |
| w/o QATC (TF-IDF-based ER) | |||||||
| TF-IDF | InfoXLM | 76.57 | 83.26 | 90.15 | 74.89 | 79.36 | 76.61 |
| XLM-R | 76.47 | 82.93 | 90.15 | 76.39 | 81.41 | 78.58 | |
| Ernie-M | 75.75 | 81.97 | 90.15 | 78.71 | 82.28 | 80.65 | |
Table 4 presents the ablation results to evaluate the contribution of each component in the SemViQA framework. When employing the full model with both the QATC-based evidence retrieval and the two-step verdict classification (TVC), SemViQA achieves the best performance across both datasets. Notably, using Ernie-Mlarge yields the highest strict accuracy (78.97%) and evidence retrieval accuracy (80.91%) on the ISE-DSC01 dataset. Removing the binary classification stage (i.e., using only one-step VC) leads to a noticeable performance drop, especially on ISE-DSC01, indicating that the binary classifier enhances the distinction between Supported and Refuted labels. Furthermore, replacing the QATC module with a TF-IDF based retriever results in a significant decline (about 5%) in evidence retrieval accuracy, which subsequently affects the overall performance. These findings highlight the critical role of both QATC and the two-step classification scheme in improving SemViQA’s effectiveness on fact verification in Vietnamese.
C.2 Analysis of Confidence Threshold in SemViQA
The confidence threshold plays a crucial role in balancing accuracy and inference time in SemViQA’s evidence retrieval process. Analysis from Figure 6 indicates that as the threshold increases from 0.0 to 0.5, evidence retrieval accuracy improves significantly, particularly on ViWikiFC ( 95%) and ISE-DSC01 ( 80.8%). However, beyond 0.5, accuracy gains plateau, while inference time decreases sharply due to the system filtering out low-confidence evidence more aggressively. Setting an optimal threshold in the range of 0.4 - 0.5 achieves a trade-off between efficiency and accuracy, ensuring that SemViQA operates swiftly while maintaining precise evidence retrieval.
C.3 Effectiveness of QATC over Traditional QA-based Evidence Retrieval
| SemViQA | ViWikiFC | ISE-DSC01 | |||||
| SER | TVC | Strict Acc | VC Acc | ER Acc | Strict Acc | VC Acc | ER Acc |
| QATC-based ER | |||||||
| InfoXLM | InfoXLM | 80.68 | 83.98 | 95.31 | 75.13 | 79.60 | 76.87 |
| XLM-R | 80.82 | 83.88 | 95.31 | 76.74 | 81.71 | 78.95 | |
| Ernie-M | 80.06 | 83.17 | 95.31 | 78.97 | 82.49 | 80.91 | |
| ViMRClarge | InfoXLMlarge | 80.25 | 83.84 | 94.69 | 75.13 | 79.54 | 76.87 |
| XLM-Rlarge | 80.34 | 83.64 | 94.69 | 76.71 | 81.65 | 78.91 | |
| Ernie-Mlarge | 79.53 | 82.97 | 94.69 | 78.97 | 82.54 | 80.91 | |
| QA-based ER | |||||||
| InfoXLM | InfoXLM | 79.96 | 83.50 | 94.45 | 74.02 | 78.95 | 75.83 |
| XLM-R | 80.11 | 83.60 | 94.45 | 75.61 | 80.95 | 77.91 | |
| Ernie-M | 79.24 | 82.74 | 94.45 | 77.82 | 81.76 | 79.82 | |
| ViMRClarge | InfoXLMlarge | 79.77 | 83.84 | 94.26 | 74.05 | 78.93 | 75.87 |
| XLM-R | 79.87 | 83.79 | 94.26 | 75.65 | 80.93 | 77.95 | |
| Ernie-M | 79.01 | 82.78 | 94.26 | 77.84 | 81.73 | 79.86 | |
To compare the learning capabilities of QATC with traditional QA models, we construct two versions of the SemViQA pipeline that are identical in structure, differing only in the evidence retrieval model. As shown in Table 5, using QATC consistently outperforms QA-based retrieval across both datasets. Specifically, on ViWikiFC, QATC achieves up to 80.82% Strict Accuracy and 95.31% ER Accuracy, while QA-based models peak at 80.11% and 94.45%, respectively. The improvement is even more evident on ISE-DSC01, where QATC reaches 78.97% Strict Accuracy and 80.91% ER Accuracy. These results confirm that QATC is more effective at learning to identify relevant evidence, leading to superior performance across the entire fact-checking pipeline.
C.4 Analysis of Confusion Matrix in Verdict Classification
To evaluate the effectiveness of our Two-step Verdict Classification (TVC) strategy, we analyze the confusion matrices in Figure 7, which compare the standard three-class classification with our proposed two-step approach, focusing solely on Supported and Refuted claims. The three-class classifier achieves strong performance, with accuracy of 95.4% for Supported and 94.6% for Refuted; however, it still exhibits notable confusion between the two classes, likely due to their semantic proximity and shared evidence patterns. In contrast, the two-step approach—by isolating NEI cases early and applying a dedicated binary classifier further improves accuracy to 98.0% for Supported and 96.4% for Refuted. These results support our hypothesis that decomposing the verdict classification task into semantically coherent subtasks enhances the model’s precision in detecting factual consistency.
Appendix D Hyperparameter and LLM Training Configuration
@@
In this section, we present the detailed hyperparameter settings and training configurations for both our SemViQA models and the Large Language Model (LLM) fine-tuning process. Table 6 consolidates all hyperparameters used across different models, including Binary Classification (BC), Three-Class Classification (TC), Question Answering with Token Classification (QATC), and LLM fine-tuning.
@@
We fine-tune a Large Language Model (LLM) using a restructured version of the original datasets, ViWikiFC and ISE-DSC01, as detailed in Figure 7. These datasets have been carefully adapted for training to improve performance and ensure compatibility with our model. For training, we utilize the official Qwen LLM implementation from the QwenLM repository666https://github.com/QwenLM/Qwen. Our training setup follows the full configuration outlined in Table 6, ensuring optimal efficiency and alignment with best practices.
@@
- Remember, the evidence must be a full sentence, not part of a sentence or less than one sentence. Given a claim and context as follows: Context: The actress revealed her secrets to maintaining a youthful appearance as follows: Eating three balanced meals a day. For dinner, Ivy Chen usually eats early to ensure her body has enough time to digest food, metabolize energy, and avoid putting pressure on the stomach and other organs. A recent study published in *Frontiers in Nutrition* suggests that eating dinner earlier can lead to a longer lifespan, with the ideal time being 7 PM. If this is not possible, experts recommend having the last meal of the day 2-3 hours before bedtime. Drinking ginger tea: To keep her body warm, promote blood circulation, and enhance circulation, Ivy Chen drinks ginger tea daily. Her ginger tea is typically made with ground ginger, black tea, turmeric powder, and brown sugar. This drink is a natural remedy that not only boosts the immune system and reduces inflammation but also fights oxidation, supports weight loss, improves skin health, and helps maintain a youthful look. Regular exercise: Ivy Chen is a fitness enthusiast who loves physical activities and exercises daily, even during pregnancy. The Taiwanese actress shared that if she is not busy with work, she runs for at least 30 minutes every day. Even when traveling abroad, she maintains her running habit. A recent study published in *Progress in Cardiovascular Disease* found that regular runners live three years longer than non-runners. Running significantly helps with weight loss, maintaining a balanced physique, toning muscles, relaxing the mind, and benefiting heart health. Besides running, Ivy Chen also swims, practices yoga, and hikes to maintain physical fitness and endurance. Skincare: Regarding her skincare routine, the actress emphasized the importance of hydration. The Taiwanese beauty revealed that she always carries a facial mist to ensure her skin stays hydrated while outdoors. Claim: Even when traveling abroad, Ivy Chen maintains her running habit. Answer: This claim is classified as Supported. The evidence is: Even when traveling abroad, she maintains her running habit.
@@
Appendix E Comparison of TF-IDF and QATC in Fact-Checking: Examples of Incorrect vs. Correct Evidence Selection